Determine whether the given quadratic function has a maximum value or minimum... Then find the coordinates of the maximum or minimum point.
f(x)= x^2+2x-9

and

f(x)=-x^2-2x-6

Answer by lwsshak3(11628)

Determine whether the given quadratic function has a maximum value or minimum... Then find the coordinates of the maximum or minimum point.
f(x)= x^2+2x-9
complete the square:
f(x)= (x^2+2x+1)-1-9
f(x)= (x+1)^2-10
This is a parabola that opens upwards with a minimum of -10 at x=-1
coordinates:(-1,-10)
and
f(x)=-x^2-2x-6
f(x)=-(x^2+2x+1)+1-6
f(x)=-(x+1)^2-5
This is a parabola that opens downwards with a maximum of -5 at x=-1
coordinates:(-1,-5)